The balance of a saving account compounded continuously will increase by 70% in 12 years. Find the annual interest rate of the account
Answer by lwsshak3(11628)

The balance of a saving account compounded continuously will increase by 70% in 12 years. Find the annual interest rate of the account.
***
Formula for continuous compounding: A=Pe^rt, P=initial investment, r=annual interest rate, t=period in years, A=amt after t years
For given problem:
t=12
A/P=e^rt
1.7=e^rt
take log of both sides
ln(1.7)=rt*lne
lne=1
rt=ln(1.7)
r=ln(1.7)/t=ln(1.7)/12≈4.42%
annual interest rate of the account=4.42%